OpenStudy (anonymous):

Integral of 1/ sqrt(1+2x) dx ? Upper limit=12 Lower limit = 0.

OpenStudy (anonymous):

1/5 - 1 = -4/5

OpenStudy (anonymous):

\[ \int\limits_{0}^{12} 1 \div \sqrt{1+2x}\] this is what it looks like. im getting 8 but the answer is 4 ..

OpenStudy (anonymous):

Work: Upper limit - 1/sqrt(1 + 24) = 1/sqrt(25) = 1/5 Lower limit - 1/sqrt(1 + 0) = 1/1 = 1 Upper limit - lower limit = 1/5 - 1 = -4/5

OpenStudy (anonymous):

but the answer is 4 ..

OpenStudy (anonymous):

Oh, crap... I forgot how to integrate, lawl. Let's try that again...

OpenStudy (anonymous):

1 + 2x = u 2dx = du dx = du/2 1/2 * integral(1/rad(u) du) I think this is it.

OpenStudy (anonymous):

Hmm, okay, thanks..